OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x - y = 3 zero −3 3 2

OpenStudy (anonymous):

First multiply second equation by 2. Add the first equation and the new equation formed to get x..

OpenStudy (anonymous):

Got it... ? o.O

OpenStudy (anonymous):

2x-2y=6?

OpenStudy (anonymous):

Yeah..Now add the equations...

OpenStudy (anonymous):

8x + 4y = 18?

OpenStudy (anonymous):

Like this..\[4x+2y+(2x-2y)=12+6\]

OpenStudy (anonymous):

Read this before answering.. http://openstudy.com/code-of-conduct @Varuns

OpenStudy (anonymous):

Hadn't seen that before. Thanks :) You can eliminate y by multiplying the second equation by -2 and then subtracting Eqn 2 from Eqn 1

OpenStudy (anonymous):

@Jessica_Moore Try again.... Never mind.. ;) @Varuns

OpenStudy (anonymous):

I got it now, thanks!!